OpenStudy (anonymous):

A new car that sells for $19,000 depreciates (decreases in value) 21% each year. What is the decay factor, b? 79 0.79 0.21 21

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

whats your guess?

OpenStudy (anonymous):

0.21

OpenStudy (anonymous):

look at you bein all smart ms smarty pants lol

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

oml lol YES :P

OpenStudy (anonymous):

How is the graph of log (x – 2) + 1 translated from the graph of log x? shifted left 1 unit and down 2 units shifted right 2 units and up 1 unit shifted right 1 unit and down 2 units shifted left 2 units and up 1 unit

OpenStudy (anonymous):

youre asking the wrong person about logarithms lol. just being honest

OpenStudy (anonymous):

i dont remember how to do them really but ill try

OpenStudy (anonymous):

I got it nvm!

OpenStudy (anonymous):

good good. i would think it would be D

OpenStudy (anonymous):

An investment of $450 increases at a rate of 6.5% per year. What is the growth factor, b? 0.65 65 0.065 6.5

OpenStudy (anonymous):

what is your guess?

OpenStudy (anonymous):

0.065

OpenStudy (anonymous):

wow youre on a roll

OpenStudy (anonymous):

Write an exponential function y = ab^x for a graph that includes (1, 15) and (0, 6). = 3(5)x y = 2.5(6)x y = 6(2.5)x y = 5(3)x

OpenStudy (anonymous):

are the x's supposed to be ^x's??

OpenStudy (anonymous):

if so, its C

OpenStudy (anonymous):

how is the graph of y=-1/2 (5)^x-3 translated from the graph y=1/2 (5)^x reflected across the y-axis and 3 units down reflected across the x-axis and 3 units down reflected across the x-axis and 3 units left reflected across the y-axis and 3 units right

OpenStudy (anonymous):

ok based on the graph that i made, the answer would be

OpenStudy (anonymous):

B. reflected across the x-axis and 3 units down

OpenStudy (anonymous):

hell yeah i passed the test!!